OpenStudy (acespeedfighter):

What is the value of 1/2 x3 + 3.4y when x = 4 and y = 2? (4 points) 12.8 14.8 25.8 38.8

OpenStudy (sachintha):

\(x^3 = x^2\times x= x\times x\times x\) Now try to write the expression by substituting the values.

OpenStudy (acespeedfighter):

hmmm

OpenStudy (acespeedfighter):

so its C

OpenStudy (sachintha):

Nope. Show me the expression after substituting the values.

OpenStudy (acespeedfighter):

Im not good at math sadly ;-;

OpenStudy (sachintha):

\(\frac{1}{2}x^3+3.4y=\frac{1}{2}(4)^3+3.4(2)\)

OpenStudy (acespeedfighter):

idk ;-;

OpenStudy (acespeedfighter):

brb

OpenStudy (prizzyjade):

\[\frac{ 1 }{ 2 } x^{3}+ 3.4y \] substitute then the given value of y and x

OpenStudy (sachintha):

\(\frac{1}{2}x^3+3.4y=\frac{1}{2}(4\times 4\times 4)+3.4\times2\)

OpenStudy (prizzyjade):

\[\frac{ 1 }{ 2 }x ^{3} + 3.4y = \frac{ 1 }{ 2 }(4^{3})+3.4(2)\]

OpenStudy (prizzyjade):

since \[4^{3}=4\times4\times4=16\times4=64\] then \[\frac{1}{2} (4^3) + 3.4(2) = \frac {1}{2} (64)+ 3.4(2)\] we can see that \[\frac {1}{2} (64) = \frac {64}{2}\] which is equal to 32. \[32 + 3.4(2)= 32+6.8\] \[=38.8\]